A subway train starting from rest leaves a
station with a constant acceleration. At the
end of 7.98 s, it is moving at 15.2418 m/s. What is the train’s displacement in the ?rst 5.19498 s of motion? Answer in units of m


 



Grade:11

3 Answers

Nishant Chandrashekar
10 years ago

v=u + at

where u=0    (since it starts from rest)

from this,

a=(15.2418)/7.98

now use ,

s=ut+(1/2)at^2

where u=0

s=25.7m

hardik joshi
10 years ago

hey,guy,is there two que.?

 

 

 

 

if there is 2 questions,,,

so,the first ans is 315.72m

hardik joshi
10 years ago

HEY,FRIEND,,

IS THERE TWO QUE.?

 

 

IF,IT IS,,,,

SO,THE FIRST ANS  WOULD BE 485.30m